# Bridging Cache-Friendliness and Concurrency: A Locality-Optimized In-Memory B-Skiplist

### v2

@@ Abstract.
On 128128128 threads, the proposed concurrent B-skiplist achieves between 2×2$\times$2 ×9×9$\times$9 × higher throughput compared to state-of-the-art concurrent skiplist implementations, including Facebook’s concurrent skiplist from Folly and the Java ConcurrentSkipListMap. Furthermore, we find that the B-skiplist achieves competitive (0.9×0.9$\times$0.9 ×1.7×1.7$\times$1.7 ×) throughput on point workloads compared to state-of-the-art cache-optimized tree-based indices (e.g., Masstree). For a more complete picture of the performance, we also measure the latency of skiplist- and tree-based indices and find that the B-skiplist achieves between 3.5×3.5$\times$3.5 ×103×103$\times$103 × lower 99% latency compared to other concurrent skiplists and between 0.85×0.85$\times$0.85 ×64×64$\times$64 × lower 99% latency compared to tree-based indices on point workloads with inserts.
@@ 1. Introduction
Table 2 shows that skiplists incur significantly more cache misses than B-trees, leaving performance on the table. Concretely, on the tested workloads, a state-of-the-art skiplist (fol, [n. d.]) incurs 2.44.8×2.4-4.8$\times$2.4 - 4.8 × more cache misses than a comparable B-tree (Xu et al., 2023) on both point and range workloads. This discrepancy in cache misses translates into actual performance: the B-tree achieves between 2×2$\times$2 ×-8×8$\times$8 × higher throughput than the skiplist on these workloads.
@@
Theoretically, given some node size $B=\Theta(Z)$, the straightforward way to improve the cache-friendliness of skiplists is to promote elements with probability $1/B$ rather than 1/21/21 / 2 as in traditional skiplists. Indeed, Golovin (Golovin, 2010) proposed this method with the B-skiplist, a cache-optimized skiplist that matches the B-tree’s bounds in expectation. For each level in this theoretical B-skiplist, consecutive unpromoted elements are stored in the same node. However, this initial paper on B-skiplists stops short of providing theoretical guarantees w.h.p., parallelization, and implementation, leaving a massive gap in making B-skiplists practical.
@@
  • An empirical evaluation of the concurrent B-skiplist compared to several in-memory skiplist-based and B-tree-based indexes that demonstrates that the B-skiplist achieves between 2×2$\times$2 ×9×9$\times$9 × higher throughput on YCSB workloads (Cooper et al., 2010) compared to state-of-the-art concurrent skiplists.
Load A B C E 111555101010Workload Normalized throughput NHSFolly SLJava SLB-skiplist
@@
Results summary. Figure 1 demonstrates that the B-skiplist achieves between 2×2$\times$2 ×9×9$\times$9 × higher throughput on workloads from the popular Yahoo! Cloud Serving Benchmark (YCSB) (Cooper et al., 2010) compared to state-of-the-art concurrent skiplists including Facebook’s concurrent skiplist from the Folly library (fol, [n. d.]), the Java ConcurrentSkipListMap (Java, [n. d.]), and the No Hot Spot Skiplist (NHS) (Crain et al., 2013). The YCSB workloads include a mix of point operations (finds/inserts) and range operations.
Furthermore, we also evaluate two tree-based indices: a state-of-the-art concurrent B-tree (Xu et al., 2023) and Masstree (Mao et al., 2012), a popular cache-optimized in-memory index. Compared to tree-based indices, we find that the B-skiplist achieves competitive throughput (between 0.9×0.9$\times$0.9 ×1.7×1.7$\times$1.7 ×) on point workloads, and between 0.7×0.7$\times$0.7 ×7.5×7.5$\times$7.5 × throughput on range workloads.
In addition to throughput, we evaluate all data structures on latency as well for a more complete picture of data-structure performance. As we shall detail in Section 5, the B-skiplist achieves between 3.5×3.5$\times$3.5 ×103×103$\times$103 × lower 99% latency compared to existing concurrent skiplists.
@@ 2.1. Skiplist structure and operations
Structure. The skiplist (Pugh, 1990b) is a data structure that stores a hierarchy of levels of linked lists. Each linked list is sorted by the keys of the elements in that linked list, with sentinels at the beginning and end for $\infty$- ∞ and $\infty$. The bottommost level (level 0) contains all of the elements in the data structure. The list size decreases by a constant factor in expectation at each successive level as we move up the hierarchy. Each linked list at some level $\ell>0$ contains a subset of the elements in the linked list below it (level $\ell-1$). This property is called inclusion invariant, where every element present at level $\ell$ must also be present at levels $0,\ldots,\ell-1$.
@@
[…]
Operations. All operations in a skiplist start at the upper left sentinel ($\infty$- ∞) at the highest level and traverse through the pointer structure in a left-to-right and top-to-bottom fashion. The skiplist is a self-balancing data structure and does not require pointer rotation to maintain its bounds. A skiplist with $n$ elements supports all point operations (insert, delete, find) in $O(\log n)$ time in expectation and with high probability.
@@ 2.2. B-skiplist structure and operations
To find a key k in a B-skiplist, the traversal begins at a curr node initialized to the upper left sentinel (with header $\infty$- ∞), just as in the regular skiplist. The search then examines the header of curr->next. If it is less than k, the curr node is updated to curr->next. We repeat this left-to-right traversal until the header of curr->next is greater than k. At that point, we search within curr for the pred element and follow its down pointer. The search continues in this way until we reach the leaf level, at which point we determine if k is present. The main difference from a traditional skiplist is that the pred element may be in the same node as k.
@@ 3. Top-Down Insertion Algorithm
The main challenge comes when the element k is promoted to level $h>0$ because the corresponding down pointer should point to a new node (due to a split) at level $h-1$ that has not yet been created in a naive top-down traversal. Recall from Section 2 that in a B-skiplist, every header element in a node at level $i$ must have been promoted to level $i+1$. It is clear which node to set the down pointer to with a “bottom-up” algorithm, because the corresponding node has been created at level $h-1$ before the down pointer at level $h$ needs to be set. However, with the proposed “top-down” algorithm, the node with the new element k has not yet been created at level $h-1$ since we are going from $h$ down to 0.
@@
To determine where to splice in the other new nodes in levels 0 to $h-1$, we continue the traversal level-by-level. Let us consider the levels in turn, starting with level $h-1$. Just as in the higher levels, the traversal will follow the down pointer corresponding to the prev element in level $h$ which will point to some node at level $h-1$. At level $h-1$, the traversal will proceed left-to-right until we find some node x such that x->header < k but x->next->header > k. Let n be the preallocated node destined to be at level $h-1$. Elements greater than k in node x should be moved to n. It is then linked into the B-skiplist by setting n->next = x->next and x->next = n. The insertion then proceeds in this way until it reaches the leaf level by following the down pointer of the last element in x, which is the largest element at that level less than k. The element k will appear in the correct nodes with the correct structure at the end of this top-down method with splits on the way down.
@@ 4. Top-down concurrency control
If there need to be splits at lower levels (i.e., $h>0$), we use HOH locking from top-to-bottom and left-to-right. Consider the case of a split at level $h-1$. To find the starting point of the traversal at level $h-1$, we first take the write lock on the node pointed to the by down pointer associated with the prev element in level $h$. Just as in the single-threaded case, the traversal will then proceed left-to-right until we find the two nodes that the new preallocated node will need to be inserted between. However, in the concurrent case, we take writer locks in HOH fashion left-to-right for thread safety. The actual split mechanism is unchanged from the serial case, and this process is repeated until the leaf level.
@@ 5. Evaluation
Result summary. At a high level, the B-skiplist achieves 2×2$\times$2 ×9×9$\times$9 × higher throughput and 3.5×3.5$\times$3.5 ×103×103$\times$103 × lower latency than non-cache-optimized concurrent skiplists. Furthermore, the B-skiplist achieves 0.9×0.9$\times$0.9 ×1.7×1.7$\times$1.7 × throughput on point workloads and 0.7×0.7$\times$0.7 ×7.5×7.5$\times$7.5 × throughput on range workloads compared to tree-based indices.
Systems setup. All experiments were run on a server with 64-core 2-way hyperthreaded Intel Xeon Gold 6338 CPU @ 2.00GHz with 1008 GB of memory. The server has a 3 MiB L1 data cache, a 2 MiB L1 instruction cache, a 80 MiB L2 cache and a 96 MiB L3 cache. We ran all experiments with 646464 physical cores and 128128128 hyperthreads.
@@
[…]
@@ 5.1. Comparison to skiplist-based indices
As shown in Figure 1, the B-skiplist significantly improves upon the throughput of the other skiplists on all operations. For the load phase (all inserts), B-skiplist achieves about 2×2$\times$2 ×, $\times1$, and 2.1×2.1$\times$2.1 × higher throughput, respectively, compared to the Java skiplist (JSL), the no hot spot skiplist (NHS), and the Folly skiplist (FLY). On point workloads (Workloads A, B, and C) with finds/inserts, the B-skiplist is about 4.6-6.6$\times$ faster than JSL, 5-9$\times$ faster than NHS, and 3.1-3.3$\times$ faster than FLY. For range queries (Workload E), the B-skiplist is about 9×9$\times$9 × faster than JSL and about 6×6$\times$6 × faster than NHS and FLY. All systems achieves better throughputs on the zipfian workloads than on uniform workloads, but the their relative performance to each other remains almost the same.
Additionally, as shown in Figure 6, the B-skiplist achieves lower latency across all workloads in all tested percentiles (50%, 90%, 99%, and 99.9%). The most competitive non-blocked skiplist is the one from Folly, which has at least 3.5×3.5$\times$3.5 × and often at least 3.7×3.7$\times$3.7 × higher latency compared to the B-skiplist in the different percentiles.
The B-skiplist achieves better throughput and latency compared to unblocked skiplists because it reduces cache misses with better spatial locality in the nodes, as shown in Table 2. The folly skiplist, the fastest of the state-of-the-art skiplists, incurs between 3.2-5.6$\times$ more cache misses compared to the B-skiplist. Furthermore, the B-skiplist maintains the simple structure and CC schemes that make the skiplist a popular choice for in-memory indexing.
@@ 5.2. Comparison to tree-based indices
Load A B C E 00.50.50.51111.51.51.5111Workload Normalized performance B-skiplistB-treeMassTree
@@
[…]
@@
Overall, we find that the B-skiplist achieves competitive (between 1-1.4$\times$ higher) throughput compared to the B-tree and between 1-2.1$\times$ higher throughput than Masstree on point workloads (load and A-C). We expect the B-tree and B-skiplist to be similar on read-heavy workloads because the B-skiplist has a similar height and node size compared to the B-tree. Furthermore, the B-tree only has to read one node per level, while the B-skiplist may have to take horizontal steps (following next pointers) along each level. Concretely, we found that on average, the B-skiplist takes about 1.7 horizontal steps per level in workloads A-C. As a result, the B-skiplist has a slightly lower throughput (within 0.9×0.9$\times$0.9 ×) compared to the B-tree on workload C.
The B-skiplist achieves the most consistent speedups (1.3-2.1$\times$ higher throughput) over tree-based indices on insert-heavy workloads (load and A). To understand why, we will first briefly review optimistic concurrency control (OCC) (Kung and Robinson, 1981), the CC scheme in both the B-tree and Masstree. OCC is a classical CC scheme for B-trees based on RW locks that leverages the observation that most insertions only affect the leaf level. Almost all insertions under OCC make one root-to-leaf pass with reader locks on the internal nodes. However, if an element must be promoted, the insert retires back to the root, taking write locks on all nodes on the way down. In contrast, the top-down CC scheme in the B-skiplist is guaranteed to always make one pass from root-to-leaf and to take write locks only on the levels that the element is promoted to. To measure this difference, we counted the number of times the root lock was taken in write mode (blocking all other operations) during the load phase and workload A in the B+-tree and B-skiplist. In the load phase, the B+-tree root write lock was taken 26K times, compared to 7 times in the B-skiplist. In workload A, the B-tree took the write lock on the root about 8.3K times, compared to 3 times in the B-skiplist. The B+-tree also exhibits higher latency in the 99th percentile compared to the B-skiplist due to the B+-tree’s retires back to the root.
50909999.9505050100100100150150150200200200Percentile Latency ($\mu$s) B-skiplistB-treeMassTree
@@
On the other hand, the B-tree achieves about 1.4×1.4$\times$1.4 × higher throughput than the B-skiplist on range queries (Workload E). As mentioned earlier, although the B-skiplist has $\Theta(B)$ elements per node in expectation, the number of elements actually in a logical node can vary by up to a factor of $O(\log n)$. Therefore, the average density (number of elements in a node) is lower in the B-skiplist compared to the B+-tree, which deterministically splits nodes when they become full. To measure this difference, we counted the number of nodes at the bottom level traversed in both the B-skiplist and B+-tree during workload E. On average, the B-skiplist accesses about 222 nodes per range query while the B+-tree accesses only about 1.51.51.5 nodes per range query. We note that in-memory indexes are traditionally optimized for point (OLTP) workloads and range queries are often an optional function. Future work involves improving range queries in B-skiplists by improving the average node density.
@@ 5.3. Strong Scaling
On the write-heavy workload A, all systems except Masstree scale number of threads increases. Masstree’s performance peaks at 32 threads and declines afterward. On 128 threads, NHS achieves about 26$\times$ speedup. The B-skiplist, B-tree, and Folly skiplist all achieve about 35-38$\times$ speedup, while JSL achieves about 45×45$\times$45 × speedup. Although JSL achieves higher parallel scalability, its overall throughput is lower than the C++-based data structures because it does not employ blocking. In contrast, on the read-heavy workload C, all systems achieve higher speedup compared to workload A because there are no writes and therefore less lock contention. On workload C, all systems achieve between 50-60$\times$ speedup on 128 threads except for NHS which achieves about 35$\times$ speedup.
1248163264128101010202020303030404040505050Num. ThreadsSpeedupB-skiplistFollyJSLB-treeMasstreeNoHotSpot
@@
1248163264128101010202020303030404040505050606060707070Num. ThreadsSpeedupB-skiplistFollyJSLB-treeMasstreeNoHotSpot
@@ 6. Conclusion
The empirical evaluation demonstrates that the B-skiplist achieves between 2×2$\times$2 ×9×9$\times$9 × higher throughput and between 3.5×3.5$\times$3.5 ×103×103$\times$103 × lower 99th percentile latency compared to popular state-of-the-art concurrent skiplist implementations such as those from Facebook’s folly library and the Java concurrent skiplist library. These results suggest that the B-skiplist is a good candidate for high-performance in-memory indexing because it resolves locality issues in skiplists while minimizing CC overhead.

### v3

@@ Abstract.
On $128$ threads, the proposed concurrent B-skiplist achieves between $2\times$$9\times$ higher throughput compared to state-of-the-art concurrent skiplist implementations, including Facebook’s concurrent skiplist from Folly and the Java ConcurrentSkipListMap. Furthermore, we find that the B-skiplist achieves competitive ($0.9\times$$1.7\times$) throughput on point workloads compared to state-of-the-art cache-optimized tree-based indices (e.g., Masstree). For a more complete picture of the performance, we also measure the latency of skiplist- and tree-based indices and find that the B-skiplist achieves between $3.5\times$$103\times$ lower 99% latency compared to other concurrent skiplists and between $0.85\times$$64\times$ lower 99% latency compared to tree-based indices on point workloads with inserts.
@@ 1. Introduction
Table 2 shows that skiplists incur significantly more cache misses than B-trees, leaving performance on the table. Concretely, on the tested workloads, a state-of-the-art skiplist (fol, [n. d.]) incurs $2.4-4.8\times$ more cache misses than a comparable B-tree (Xu et al., 2023) on both point and range workloads. This discrepancy in cache misses translates into actual performance: the B-tree achieves between $2\times$-$8\times$ higher throughput than the skiplist on these workloads.
@@
Theoretically, given some node size $B=\Theta(Z)$, the straightforward way to improve the cache-friendliness of skiplists is to promote elements with probability $1/B$ rather than $1/2$ as in traditional skiplists. Indeed, Golovin (Golovin, 2010) proposed this method with the B-skiplist, a cache-optimized skiplist that matches the B-tree’s bounds in expectation. For each level in this theoretical B-skiplist, consecutive unpromoted elements are stored in the same node. However, this initial paper on B-skiplists stops short of providing theoretical guarantees w.h.p., parallelization, and implementation, leaving a massive gap in making B-skiplists practical.
@@
  • An empirical evaluation of the concurrent B-skiplist compared to several in-memory skiplist-based and B-tree-based indexes that demonstrates that the B-skiplist achieves between $2\times$$9\times$ higher throughput on YCSB workloads (Cooper et al., 2010) compared to state-of-the-art concurrent skiplists.
Load A B C E 11551010Workload Normalized throughput NHSFolly SLJava SLB-skiplist
@@
Results summary. Figure 1 demonstrates that the B-skiplist achieves between $2\times$$9\times$ higher throughput on workloads from the popular Yahoo! Cloud Serving Benchmark (YCSB) (Cooper et al., 2010) compared to state-of-the-art concurrent skiplists including Facebook’s concurrent skiplist from the Folly library (fol, [n. d.]), the Java ConcurrentSkipListMap (Java, [n. d.]), and the No Hot Spot Skiplist (NHS) (Crain et al., 2013). The YCSB workloads include a mix of point operations (finds/inserts) and range operations.
Furthermore, we also evaluate two tree-based indices: a state-of-the-art concurrent B-tree (Xu et al., 2023) and Masstree (Mao et al., 2012), a popular cache-optimized in-memory index. Compared to tree-based indices, we find that the B-skiplist achieves competitive throughput (between $0.9\times$$1.7\times$) on point workloads, and between $0.7\times$$7.5\times$ throughput on range workloads.
In addition to throughput, we evaluate all data structures on latency as well for a more complete picture of data-structure performance. As we shall detail in Section 5, the B-skiplist achieves between $3.5\times$$103\times$ lower 99% latency compared to existing concurrent skiplists.
@@ 2.1. Skiplist structure and operations
Structure. The skiplist (Pugh, 1990b) is a data structure that stores a hierarchy of levels of linked lists. Each linked list is sorted by the keys of the elements in that linked list, with sentinels at the beginning and end for -$\infty$ and $\infty$. The bottommost level (level 0) contains all of the elements in the data structure. The list size decreases by a constant factor in expectation at each successive level as we move up the hierarchy. Each linked list at some level $\ell>0$ contains a subset of the elements in the linked list below it (level $\ell-1$). This property is called inclusion invariant, where every element present at level $\ell$ must also be present at levels $0,\ldots,\ell-1$.
@@
An element that appears at some maximum level $\ell>0$ is said to be promoted to that level. Promotions are determined upfront at the start of an insertion via randomization with a series of coin flips. Notably, the level that an element is promoted to in a skiplist is unrelated to the current structure of elements in the skiplist. That is, the highest level that an element appears at in a skiplist is equal to the number of successive “heads” seen when flipping a coin with some constant probabilty $p$ (usually $p=1/2$, but any probabilty $1/c$ where $c$ is a constant sufficies to achieve the asymptotic bounds). Given a skiplist with $n$ elements, the maximum height of any element is $O(\log n)$ in expectation and with high probability.
Operations. All operations in a skiplist start at the upper left sentinel (-$\infty$) at the highest level and traverse through the pointer structure in a left-to-right and top-to-bottom fashion. The skiplist is a self-balancing data structure and does not require pointer rotation to maintain its bounds. A skiplist with $n$ elements supports all point operations (insert, delete, find) in $O(\log n)$ time in expectation and with high probability.
@@ 2.2. B-skiplist structure and operations
To find a key k in a B-skiplist, the traversal begins at a curr node initialized to the upper left sentinel (with header -$\infty$), just as in the regular skiplist. The search then examines the header of curr->next. If it is less than k, the curr node is updated to curr->next. We repeat this left-to-right traversal until the header of curr->next is greater than k. At that point, we search within curr for the pred element and follow its down pointer. The search continues in this way until we reach the leaf level, at which point we determine if k is present. The main difference from a traditional skiplist is that the pred element may be in the same node as k.
@@ 3. Top-Down Insertion Algorithm
The main challenge comes when the element k is promoted to level $h>0$ because the corresponding down pointer should point to a new node (due to a split) at level h1h-1 that has not yet been created in a naive top-down traversal. Recall from Section 2 that in a B-skiplist, every header element in a node at level $i$ must have been promoted to level $i+1$. It is clear which node to set the down pointer to with a “bottom-up” algorithm, because the corresponding node has been created at level h1h-1 before the down pointer at level $h$ needs to be set. However, with the proposed “top-down” algorithm, the node with the new element k has not yet been created at level h1h-1 since we are going from $h$ down to 0.
@@
To determine where to splice in the other new nodes in levels 0 to h1h-1, we continue the traversal level-by-level. Let us consider the levels in turn, starting with level h1h-1. Just as in the higher levels, the traversal will follow the down pointer corresponding to the prev element in level $h$ which will point to some node at level h1h-1. At level h1h-1, the traversal will proceed left-to-right until we find some node x such that x->header < k but x->next->header > k. Let n be the preallocated node destined to be at level h1h-1. Elements greater than k in node x should be moved to n. It is then linked into the B-skiplist by setting n->next = x->next and x->next = n. The insertion then proceeds in this way until it reaches the leaf level by following the down pointer of the last element in x, which is the largest element at that level less than k. The element k will appear in the correct nodes with the correct structure at the end of this top-down method with splits on the way down.
@@ 4. Top-down concurrency control
If there need to be splits at lower levels (i.e., $h>0$), we use HOH locking from top-to-bottom and left-to-right. Consider the case of a split at level h1h-1. To find the starting point of the traversal at level h1h-1, we first take the write lock on the node pointed to the by down pointer associated with the prev element in level $h$. Just as in the single-threaded case, the traversal will then proceed left-to-right until we find the two nodes that the new preallocated node will need to be inserted between. However, in the concurrent case, we take writer locks in HOH fashion left-to-right for thread safety. The actual split mechanism is unchanged from the serial case, and this process is repeated until the leaf level.
@@ 5. Evaluation
Result summary. At a high level, the B-skiplist achieves $2\times$$9\times$ higher throughput and $3.5\times$$103\times$ lower latency than non-cache-optimized concurrent skiplists. Furthermore, the B-skiplist achieves $0.9\times$$1.7\times$ throughput on point workloads and $0.7\times$$7.5\times$ throughput on range workloads compared to tree-based indices.
Systems setup. All experiments were run on a server with 64-core 2-way hyperthreaded Intel Xeon Gold 6338 CPU @ 2.00GHz with 1008 GB of memory. The server has a 3 MiB L1 data cache, a 2 MiB L1 instruction cache, a 80 MiB L2 cache and a 96 MiB L3 cache. We ran all experiments with $64$ physical cores and $128$ hyperthreads.
@@
We evaluate each workload under both uniform random and zipfian distributions in the run phase. In the uniform workload, the elements in both the load and run phases are generated from a uniform distribution. In the zipfian workload, the elements in the load phase are generated from a uniform distribution, while the elements in the run phase are generated from a zipfian distribution. We omit the results on the zipfian distirbution due to space limitations, but the results were similar (on average within $20\%$) to the uniform distribution.
@@ 5.1. Comparison to skiplist-based indices
[…]
Additionally, as shown in Figure 6, the B-skiplist achieves lower latency across all workloads in all tested percentiles (50%, 90%, 99%, and 99.9%). The most competitive non-blocked skiplist is the one from Folly, which has at least $3.5\times$ and often at least $3.7\times$ higher latency compared to the B-skiplist in the different percentiles.
[…]
@@ 5.2. Comparison to tree-based indices
Load A B C E 00.50.5111.51.511Workload Normalized performance B-skiplistB-treeMassTree
@@
Systems setup. We compare the B-skiplist with a high-performance concurrent B+-tree 101010 https://github.com/wheatman/BP-Tree/tree/main/tlx-plain/container (Xu et al., 2023) (a common B-tree variant) based on optimistic concurrency control (Kung and Robinson, 1981). The default configuration for the B+-tree sets $\texttt{node\_size}=1024$ bytes. Both the concurrent B-skiplist and concurrent B+-tree use the same RW lock.
@@
[…]
@@
On the other hand, the B-tree achieves about $1.4\times$ higher throughput than the B-skiplist on range queries (Workload E). As mentioned earlier, although the B-skiplist has $\Theta(B)$ elements per node in expectation, the number of elements actually in a logical node can vary by up to a factor of $O(\log n)$. Therefore, the average density (number of elements in a node) is lower in the B-skiplist compared to the B+-tree, which deterministically splits nodes when they become full. To measure this difference, we counted the number of nodes at the bottom level traversed in both the B-skiplist and B+-tree during workload E. On average, the B-skiplist accesses about $2$ nodes per range query while the B+-tree accesses only about $1.5$ nodes per range query. We note that in-memory indexes are traditionally optimized for point (OLTP) workloads and range queries are often an optional function. Future work involves improving range queries in B-skiplists by improving the average node density.
@@ 5.3. Strong Scaling
[…]
124816326412810102020303040405050Num. ThreadsSpeedupB-skiplistFollyJSLB-treeMasstreeNoHotSpot
@@
12481632641281010202030304040505060607070Num. ThreadsSpeedupB-skiplistFollyJSLB-treeMasstreeNoHotSpot
@@ 6. Conclusion
The empirical evaluation demonstrates that the B-skiplist achieves between $2\times$$9\times$ higher throughput and between $3.5\times$$103\times$ lower 99th percentile latency compared to popular state-of-the-art concurrent skiplist implementations such as those from Facebook’s folly library and the Java concurrent skiplist library. These results suggest that the B-skiplist is a good candidate for high-performance in-memory indexing because it resolves locality issues in skiplists while minimizing CC overhead.